# Position: LLM Unlearning Benchmarks are Weak Measures of Progress

### v1

@@ I Introduction
First, we identify that unlearning algorithms are particularly vulnerable in the face of queries that have dependencies on both the unlearned data and retained data. Popular existing benchmarks typically evaluate on disjoint forget and retain sets (see e.g.  [17], Figure 1), where forget queries are related only to the unlearned data and retain queries are largely unrelated to the unlearned data. As a result, we find that simple modifications that introduce even rough dependencies between the forget and retain sets can cause catastrophic failures in unlearning methods optimized on these benchmarks.
@@ III-B Threat model
In Section IV, show that even under this extremely weak threat model, we can craft simple queries that reveal “unlearned” information or destroy the “retain” performance (utility) of the model after unlearning.
@@ IV-A TOFU: Asking about both forget and retain entities
As a simple example, we first explore the effect of introducing dependencies between forget and retain sets in the TOFU benchmark  [10]. The questions in TOFU each pertain to a single author either from the forget or retain set. We make a straightforward modification to the test queries by combining multiple questions together, asking about two authors: one from the forget set and one from the retain set. For example, we create the paired question:
@@ 2. What is the full name of the author born on July 17, 1953 in Accra, Ghana?
We quantify these results in Figure 3. We test three methods: gradient ascent, preference optimization, and ECO  [^1] We train gradient ascent and preference optimization with 1 epoch of unlearning each, with the default parameters given by the TOFU authors [10].. Our goal is to test whether the retain efficacy is the same whether a retain set query is asked independently, or paired with a forget query. For each method, we generate two sets of responses to retain queries: first, we ask the retain query alone, and second, we ask a combined query pairing a retain set query with a forget set query (using the 10% forgetting corpus). From the combined responses, we extract only the retain query response from the model where possible (if the model does not give a structured, paired response, we treat the entire answer as the retain answer). We then compute the ROUGE metrics for each set of responses by comparing it to the correct ground truth answer, expecting that better retain efficacy corresponds to higher similarity to ground truth.
@@ IV-B WMDP: Modifying multiple-choice responses
We test WMDP unlearning methods by performing a simple modification on the retain set: we randomly replace one random incorrect answer choice with the term “SARS-CoV-2”, which appears frequently in the unlearning data. Because the swapped choice is not relevant to answering the benign, retain-set query, this change should have no effect on retain performance. We test two unlearning methods: RMU, the main baseline proposed in the WMDP benchmark paper [5], and RMU with targeted latent adversarial training (LAT) [62]. Both works report MMLU (retain) performance close to the performance of the base model (Zephyr-7B) with no unlearning.
@@
The results of this experiment are in Figure 4. The key takeaway is that making this benign modification degrades performance of the unlearned model on the retain set to near-random. As a control, we also evaluate the base model (with no unlearning) on this modified dataset. We find that even the base model experiences some accuracy loss with this change (consistent with prior work investigating the general brittleness of LLMs on multiple-choice queries  [15, 11])—but much less than the unlearned model.
RMU suffers significantly more than the base model on college biology (a 2% drop for the base model vs. a 20% drop for RMU) and overall accuracy across MMLU tasks (a 12% drop for the base model vs. a 27% drop for RMU), indicating that the unlearning method itself made the model more vulnerable to this otherwise benign change in the query set. The drop in performance is similar for college chemistry, and performance of the unlearned model increases slightly for virology, but we note that the performance of RMU was already worse than random for virology and still random after swapping choices, making this “improvement” insignificant. Finally, we find that incorporating the more recent LAT [62] to the RMU method results in an even more significant drop, with the model performing worse than random in every category.
@@ V-A Conflating the forget data and test query set
In order to train an accurate classifier for WMDP, however, ECO relies on the query set rather than the provided unlearning corpus (the method ignores the unlearning corpus entirely). Closer inspection, in fact, reveals that the classifier overfits to a spurious feature in the evaluation framework. The framework automatically prepends the following phrase to every query: “The following are multiple choice questions (with answers) about topic.” Here, topic may be, for example, “college biology” or “immunology”. Because the MMLU (retain) questions are generally topics from “college” datasets while the WMDP (forget) questions do not have the keyword “college”, the ECO classifier overfits and perfectly classifies using this spurious feature. Indeed, we find that if we remove the prepended phrase, the ECO classifier performs worse than random (37% accuracy).
@@
For completeness, in Appendix -A, we include results of our own classifier trained on a held-out set of WMDP queries. The results indicate that a correctly-trained classifier could still perform reasonably well on this benchmark, but we report results with caution given the limitations we have pointed out with forget/retain-set benchmarks as well as the fact that the classifier requires access to the query set to perform well.
@@ VI Recommendations
A tempting solution to the concerns presented in this paper is to continue to develop more sophisticated benchmarks to “patch” flaws in existing benchmarks. Unfortunately, this leads to a cat-and-mouse game that is by now familiar to researchers in security and machine learning alike. Benchmarks that attempt to partition unlearning performance into cleanly separated “forget” and “retain” query performance encourage solutions that do not model realistic queries with dependencies on both the “forget” and “retain” data. However, when such dependencies are introduced, it becomes clear that simple accuracy metrics are not sufficient to measure the unlearning achieved by an algorithm. Furthermore, a method that can perform well on a more sophisticated benchmark evaluating forget-retain dependencies might still be vulnerable to other query modifications not covered by our work. Below we discuss several recommendations for the LLM unlearning community in light of these concerns.

### v2

@@ I Introduction
First, we identify that unlearning algorithms are particularly vulnerable in the face of queries that have dependencies on both the unlearned data and retained data. Popular existing benchmarks typically evaluate on disjoint forget and retain sets (see e.g.  [17], Table 1), where forget queries are related only to the unlearned data and retain queries are largely unrelated to the unlearned data. As a result, we find that simple modifications that introduce even rough dependencies between the forget and retain sets can cause catastrophic failures in unlearning methods optimized on these benchmarks.
@@ II-C Works proposing improvements to benchmark datasets and metrics
We also note that prior work advocates for probing models for weaknesses using strong attack models, including testing whether information can be “relearned” or elicited through finetuning or careful prompting [58, 59, 60], using optimizers to bypass filters [61], or other stress-testing methods including jailbreaking or query translation [55]. In our work, however, we show that methods are not robust to benchmark perturbations even under the weakest reasonable attack model. If methods can be shown to be robust to these weaker perturbations, it makes sense to extend evaluations to include these stronger threat models as more stringent measure of robustness.
@@ III-B Threat model
In Section IV, we show that even under this extremely weak threat model, we can craft simple queries that reveal “unlearned” information or destroy the “retain” performance (utility) of the model after unlearning.
@@ IV-A TOFU: Asking about both forget and retain entities
As a simple example, we first explore the effect of introducing dependencies between forget and retain sets in the TOFU benchmark  [10]. The questions in TOFU each pertain to a single author either from the forget or retain set. We make a straightforward modification to the test queries by combining multiple questions together, asking about two authors: one from the forget set and one from the retain set. [^1] We have made these queries available for use at https://huggingface.co/datasets/forgelab/tofu-pair. For example, we create the paired question:
@@ 2. What is the full name of the author born on July 17, 1953 in Accra, Ghana?
We quantify these results in Figure 3. We test three methods: gradient ascent, preference optimization, and ECO  [^2] We train gradient ascent and preference optimization with 1 epoch of unlearning each, with the default parameters given by the TOFU authors [10].. Our goal is to test whether the retain efficacy is the same whether a retain set query is asked independently, or paired with a forget query. For each method, we generate two sets of responses to retain queries: first, we ask the retain query alone, and second, we ask a combined query pairing a retain set query with a forget set query (using the 10% forgetting corpus). From the combined responses, we extract only the retain query response from the model where possible (if the model does not give a structured, paired response, we treat the entire answer as the retain answer). We then compute the ROUGE-L score for each set of responses by comparing it to the correct ground truth answer, expecting that better retain efficacy corresponds to higher similarity to ground truth. (See Appendix -A for definitions of the ROUGE precision, recall, and F1 scores.)
@@ IV-B WMDP: Modifying multiple-choice responses
We test WMDP unlearning methods by performing a simple modification on the retain set: we randomly replace one random incorrect answer choice with the term “SARS-CoV-2”, which appears frequently in the unlearning data. Because the swapped choice is not relevant to answering the benign, retain-set query, this change should have no effect on retain performance. (We filter out any queries that include the terms “EXCEPT” or “NOT,” where “SARS-CoV-2” may be an acceptable answer choice. This filters 265 questions from the overall query set, 21 questions from college biology, 4 from college chemistry, and 0 from virology.) [^3] We have made these queries available for use at https://huggingface.co/datasets/forgelab/wmdp-swap.
We test two unlearning methods: RMU, the main baseline proposed in the WMDP benchmark paper [5], and RMU with targeted latent adversarial training (LAT) [66]. Both works report MMLU (retain) performance close to the performance of the base model (Zephyr-7B) with no unlearning.
@@
The results of this experiment are in Figure 4. The key takeaway is that making this benign modification degrades performance of the unlearned model on the retain set to near-random. As a control, we also evaluate the base model (with no unlearning) on this modified dataset. We find that the base model accuracy generally increases after swapping an incorrect choice.
In contrast, RMU suffers significantly on college biology (a nearly 30% drop in accuracy) and overall accuracy across MMLU tasks (a 28% drop in accuracy), indicating that the unlearning method itself made the model more vulnerable to this otherwise benign change in the query set. We see smaller drops in performance on college chemistry and virology (noting that performance on virology was already worse than random). Finally, we find that incorporating the more recent LAT [66] to the RMU method results in an even more significant drop, with the model performing worse than random in every category.
@@ V-A Conflating the forget data and test query set
In order to train an accurate classifier for WMDP, however, ECO [^5] This text refers to arXiv preprint arXiv:2406.07933 version 1, which was the current version at the time our work was published. relies on the query set rather than the provided unlearning corpus (the method ignores the unlearning corpus entirely). Closer inspection, in fact, reveals that the classifier overfits to a spurious feature in the evaluation framework. The framework automatically prepends the following phrase to every query: “The following are multiple choice questions (with answers) about topic.” Here, topic may be, for example, “college biology” or “immunology”. Because the MMLU (retain) questions are generally topics from “college” datasets while the WMDP (forget) questions do not have the keyword “college”, the ECO classifier overfits and perfectly classifies using this spurious feature. Indeed, we find that if we remove the prepended phrase, the ECO classifier performs worse than random (37% accuracy).
@@
For completeness, we include results of our own classifier trained on a held-out set of WMDP queries.
Method Bio Cyber Chem
Base 63.7 44.0 45.8
SSD 50.2 35.0 33.8
RMU 31.2 28.2 45.8
LC (all) 32.8 40.5 38.2
LC (holdout) 31.7 39.8 34.8
(a) WMDP Results (↓)
Method C. Bio Virology C. CS Cybersec All MT
Base 68.1 52.4 50.0 65.0 58.1 7.33
SSD 46.5 38.0 35.0 52.0 40.7 5.48
RMU 63.2 25.9 49.0 45.0 57.1 7.10
LC (all) 69.4 50.0 50.0 66.0 55.7 6.45
LC (holdout) 69.4 49.4 46.0 66.0 54.1 6.45
(b) MMLU and MT-Bench Results (↑)
TABLE II: Performance comparison of different methods on benchmarks from  [5]. Results for SSD and RMU are as reported in  [5]. We abbreviate column names for space: C. Bio and C. CS are college biology and computer science; MT refers to MT-Bench. We omit the results for LLMU and SCRUB because our results are strictly better on all columns, except for SCRUB WMDP-Bio, where SCRUB is 5% better than filtering.
We evaluated a linear classifier for the WMDP benchmark by training a binary classification head on top of a LLaMA-2-7b model. Training the classifier on the WMDP training corpora resulted in random classification performance in the test corpus (MCQ), likely because the training corpus consists of scraped articles that are very different from the MCQ format. Instead, we create a 50% train-test split of the MCQ evaluation corpora (forget data from WMDP and retain data from MMLU) and train the binary classifier on this train split. In a practical pipeline, this training data could come from retraining on real user queries after deployment.
In Table II, we report the accuracy for the linear classifier on both the held-out test set as well as the complete set of results (as we were unable to evaluate the baseline methods on only the holdout set). The linear classifier performs comparably to RMU (the given baseline) in several categories, and performs comparably to or better than RMU on the retain set.
These results show that a finetuned classifier can be a viable solution on this benchmark, but we report results with caution given the limitations we have pointed out with forget/retain-set benchmarks as well as the fact that the classifier requires access to the query set to perform well.
@@ VI Recommendations
Benchmarks that attempt to partition unlearning performance into cleanly separated “forget” and “retain” query performance encourage solutions that do not model realistic queries with dependencies on both the “forget” and “retain” data. However, when such dependencies are introduced, it becomes clear that simple accuracy metrics are not sufficient to measure the unlearning achieved by an algorithm. Furthermore, a method that can perform well on a more sophisticated benchmark evaluating forget-retain dependencies might still be vulnerable to other query modifications not covered by our work. Below we discuss several recommendations for the LLM unlearning community in light of these concerns.